General Products has an asset beta of 1.0, and a debt to equity ratio of 1.3. Their tax rate is 0.40. If the risk free rate is 0.05, and the expected return on the S&P500 is 0.18, what is the cost of levered equity for General Products?
Answer #1

Cost of levered equity=risk free rate+asset beta*(1+(1-tax rate)*Debt/Equity)*(market return-risk free rate)=0.05+1*(1+(1-0.40)*1.3)*(0.18-0.05)=0.28140000